### v6

@@ B. Assumptions on $G$:
Remark 4.3.
Assumption 4.2 (B1) is related to the convexity of function $G$ but is less stringent than the convexity condition. If it is not satisfied, similar to other gradient descent algorithms, there is a possibility for some particles to get trapped in the local minimizers $\hat{v}$ of $G$, i.e. $\nabla g(\hat{v})=0$. Nevertheless, provided the function values $\mathcal{E}(v)$ at those local minimizers of $G$ do not fall below $\mathcal{E}(v^{*})$, a condition attainable by adding a positive scalar multiple of $G$ to $\mathcal{E}$ without altering the solution $v^{*}$, it will not affect the convergence of the consensus point to the constrained minimizer $v^{*}$, as evidenced in the experiments detailed in Section 5.2.1, Figure 5. It is noteworthy that this slight adjustment on $\mathcal{E}(v)$ differs from the penalization method outlined in [7]. Here, there is no necessity for the penalty parameter to approach infinity, as the convergence is enforced through the dynamics rather than penalization. The introduction of a positive scalar multiple of $G$ to $\mathcal{E}$ is to avoid the extreme case. Thus a mild penalization would suffice.
Assumption 4.2 (B2) is primarily technical in nature. Assumption 4.2 (B3) guarantees that the gradient of $G$ around the constraint $\{G=0\}$ does not vanish too rapidly.
@@
Remark 4.4.
The above Assumption 4.2 (C1) ensures the geometry of $\mathcal{E}$ on the set $\{G=u\}$ is similar among small enough $u$, i.e., on sets $\{G=u\}$, the constrained minimizers $v_{u}$ and constrained minimums $\mathcal{E}(v_{u})$ are close among small enough values for $u$. To illustrate, if this condition is not met, as depicted in Figure 3 (a), the desired constrained minimizer $v^{*}$ (depicted as a solid green pentagon) is considerably distant from the minimizer $v_{u}$ on a nearby level set $\{G=u\}$ (depicted as a solid orange circle) for all sufficiently small $u$. Consider an extreme case where we assume $\mathcal{E}(v_{u})$ is significantly smaller than $\mathcal{E}(v^{*})$ for all positive but sufficiently small $u$. Due to the nature of gradient descent on $G$ for each particle, which may not precisely enforce each particle to remain on the constraint $\{G=0\}$, these particles will tend to remain in a neighborhood of $\{G=0\}$. Consequently, numerous particles will cluster around $\{G=u\}$ for sufficiently small $u$, as illustrated in Figure 3 (a). Given that numerous particles are near $v_{u}$, where the function value is significantly small, the algorithm computes the consensus point around $v_{u}$ rather than $v^{*}$. Consequently, the consensus point will gradually lead particles to concentrate around $v_{u}$ rather than $v^{*}$, as illustrated in Figure 3 (b), which leads to a failure in this extreme scenario. To avoid the occurrence of such extreme cases, we proposed Assumption 4.2 (C1). In conjunction with other assumptions, the similarity of the local geometry of $\mathcal{E}$ on the set $\{G=u\}$ for sufficiently small values of $u$ is guaranteed as established in Lemma 4.6.
@@
Figure 3: The blue curves represent function values on the constraint set $\{G=0\}$ and the red curves on the level set $\{G=u\}$. Dashed lines represent corresponding constraint sets. The green pentagon denotes the constrained minimizer $v^{*}$, and the orange circle represents the minimizer $v_{u}$ on the nearby level set G=u{G=u}. Empty circles represent particles.
4.3 Sketch of the Proof
@@
We plug $\tfrac{1}{2}\|v-v^{*}\|_{2}^{2}$ into Definition 3.2, which yields the following differential inequality that describes the dynamics of of the energy functional $\mathcal{V}\big{(}\rho_{t}\big{)}$.
@@
(Proof in Appendix D) Let $\mathcal{V}\big{(}\rho_{t}\big{)}$ be the energy functional defined in (9). Under Assumption 4.2,
(12) $\dfrac{d}{dt}\mathcal{V}\big{(}\rho_{t}\big{)}\leq$ $-(2\lambda-\sigma^{2})\mathcal{V}\big{(}\rho_{t}\big{)}+\sqrt{2}(\lambda+\sigma^{2})\sqrt{\mathcal{V}\big{(}\rho_{t}\big{)}}\|v_{\alpha}\big{(}\rho_{t}\big{)}-v^{*}\|_{2}+\dfrac{\sigma^{2}}{2}\|v_{\alpha}\big{(}\rho_{t}\big{)}-v^{*}\|_{2}^{2}$
$-\dfrac{1}{\epsilon}\int\Bigl{\langle}\nabla G,v-v^{*}\Bigr{\rangle}\,d\rho_{t}(v).\vskip-6.45831pt$
@@
to which Gronwall’s inequality can be applied, ensuring exponential decay. To this end, we first present the following auxiliary lemma, which draws out additional consequences from Assumption 4.2 (C).
@@
(Proof in Appendix B) There exist non-negative increasing function $\tau_{2}(x),\tau_{3}(x)$ and $\tau_{4}(x)$ mapping from $\mathbb{R}$ to $\mathbb{R}$ with $\lim_{x\rightarrow 0}\tau_{i}(x)=0$ for $i=2,3,4$ so that the following hold $\forall u,r\geq 0$ small enough:
@@
(Proof in Appendix F) Fix $r>0$ small enough and $u>0$ small enough. $q>0$ is a constant such that $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ is true $\forall\tilde{u}\in[0,u)$. Then
(14) $\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}\leq$ $2\sqrt{d}\cdot\dfrac{\big{(}q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\big{)}^{\mu}}{\eta}$
@@
It is observed that with an appropriate choice of $q,u,r$, provided $\rho_{t}(B(v^{*},r))$ is suitably bounded from below, as proven in Lemma 4.9 in Subsection 4.4, letting $\alpha$ be sufficiently large will make the first four terms above small enough. Concerning the last term of (12), it is related to $\int G\,d\rho_{t}(v)$, which can be controlled by Lemma 4.10 in Subsection 4.4. Consequently, we can control $\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$ in such a way that (13) holds, thereby ensuring exponential decay.
Remark 4.8.
The framework of the proof strategy was introduced in [22]. However, in our case, the Laplace principle shown in [22] is not directly applicable. There are two challenges in our proof. First, a global valley-like structure around $v^{*}$ is required to use the Laplace principle. In the constrained case, such a structure does not hold. But one can notice that, on a nearby level set $\{G=u\}$, if a comparable constrained valley-like structure exists for the corresponding constrained minimizer $v_{u}$, a constrained Laplace principle on $\{G=u\}$ can be established for small $u$. Integrating over all $u$ yields reasonable estimates on $\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|$. Secondly, there is always mass around the unconstrained global minimum. Consequently, increasing $\alpha$ would attract all particles towards the unconstrained global minimum rather than the constrained one. Therefore, one needs to find the balance between the gradient force and the tendency towards the consensus point.
@@
4.4 Proof of Theorem 4.1
In this subsection, we present the complete proof of Theorem 4.1. For simplicity, in the following, we assume $\tau_{1}(u)=\tau_{2}(u)=\tau_{3}(u)=\tau_{4}(u)=u$, where $\tau_{1},\ \tau_{2},\ \tau_{3},\ \tau_{4}$ are defined in Assumption 4.2 (C) and Lemma 4.6. We point out that the proof technique remains valid for any choice of $\tau_{i}$ that is an increasing function and converges to 0 as $u$ approaches 0.
@@
(Proof in Appendix G) Let $B=\sup_{t\in[0,T]}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{\infty}$. Then for all $t\in[0,T]$,
@@
Dynamics of $\int G\,d\rho_{t}(v)$ In Lemma 4.5, we have gained control over $\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$, yet the last term in the dynamics (12) remains to be studied, which we do now.
@@
(Proof in Appendix H) Assume $\sup_{t\in[0,T]}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}<\infty$ and $\sup_{t\in[0,T]}\mathcal{V}\left(\rho_{t}\right)<\infty$. Then for $\epsilon>0$ small enough, $\int G\,d\rho_{t}(v)\leq\int G\,d\rho_{0}(v)$ $\forall t\in[0,T]$.
@@
Proof 4.11 (Proof of Theorem 4.1).
First we use Lemma 4.5 to derive the dynamics of $\mathcal{V}\left(\rho_{t}\right)$:
$\dfrac{d}{dt}\mathcal{V}\big{(}\rho_{t}\big{)}\leq-(2\lambda-\sigma^{2})\mathcal{V}\big{(}\rho_{t}\big{)}+\sqrt{2}(\lambda+\sigma^{2})\sqrt{\mathcal{V}\big{(}\rho_{t}\big{)}}\|v_{\alpha}\big{(}\rho_{t}\big{)}-v^{*}\|_{2}+\dfrac{\sigma^{2}}{2}\|v_{\alpha}\big{(}\rho_{t}\big{)}-v^{*}\|^{2}_{2},$
where the last term on the right-hand-side of Lemma 4.5 is omitted because of its non-positivity due to Assumption 4.2 (B1).
Now we define $C(t)$ as $C(t)=\min\left\{\tfrac{\tau}{2}\tfrac{(2\lambda-\sigma^{2})}{\sqrt{2}(\lambda+\sigma^{2})},\sqrt{\tau\tfrac{(2\lambda-\sigma^{2})}{\sigma^{2}}}\right\}\sqrt{\mathcal{V}\left(\rho_{t}\right)}$ , and $T_{\alpha,\epsilon}$ as $T_{\alpha,\epsilon}=\sup\Big{\{}t\geq 0|\text{ }\mathcal{V}\big{(}\rho_{t^{\prime}}\big{)}>\delta,\big{\|}v_{\alpha}(\rho_{t^{\prime}}\big{)}-v^{*}\|_{2}\leq C(t^{\prime})\text{ for all $t^{\prime}\in[0,t]$}\Big{\}}.$ As long as $\|v_{\alpha}\big{(}\rho_{t^{\prime}}\big{)}-v^{*}\|_{2}\leq C(t^{\prime})$ is true, it is straightforward to verify that $\tfrac{d}{dt}\mathcal{V}\left(\rho_{t^{\prime}}\right)\leq-(1-\tau)(2\lambda-\sigma^{2})\mathcal{V}\left(\rho_{t^{\prime}}\right).$ Thus by Gronwall’s inequality, if $t\leq T_{\alpha,\epsilon}$, one has $\mathcal{V}\left(\rho_{t}\right)\leq\mathcal{V}(\rho_{0})\exp\left(-(1-\tau)(2\lambda-\sigma^{2})t\right).$
@@
One can verify that this choice of $q,r$ and $u$ will satisfy the assumptions of Lemma 4.7, i.e., $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ (for details, please see Appendix I).
@@
$\|v_{\alpha}\left(\rho_{T_{\alpha,\epsilon}}\right)-v^{*}\|_{2}\leq 2\sqrt{d}\cdot\dfrac{\big{(}q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\big{)}^{\mu}}{\eta}$
@@
(17) $2\sqrt{d}\cdot\dfrac{\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}\leq 2\sqrt{d}\cdot\dfrac{\left(4\cdot\tfrac{1}{4}\cdot\left(\eta\tfrac{C\left(T_{\alpha,\epsilon}\right)}{8\sqrt{d}}\right)^{1/\mu}\right)^{\mu}}{\eta}=\dfrac{C\left(T_{\alpha,\epsilon}\right)}{4},$
@@
where $a(\delta)=2d\max\left\{\tfrac{\lambda\left(\sqrt{c}R_{0}+C(0)\right)\sqrt{c}}{(1-c)^{2}r(\delta)}+\tfrac{\sigma^{2}\left(cR_{0}^{2}+C(0)^{2}\right)(2c+1)}{(1-c)^{4}r(\delta)^{2}},\tfrac{2\lambda^{2}}{(2c-1)\sigma^{2}}\right\}$. In the first inequality above, we used (18), the fact that $T_{\alpha,\epsilon}<T^{*}$ and Lemma 4.9 with parameter $B=\sup_{t\in[0,T_{\alpha,\epsilon}]}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{\infty}\leq\sup_{t\in[0,T_{\alpha,\epsilon}]}\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}\leq\sup_{t\in[0,T_{\alpha,\epsilon}]}C(t)\leq C(0)$. In the second inequality above, we used the Cauchy inequality. Also, Assumption 4.2 (C1) was used to deduce $\|v-v_{G(v)}\|_{2}\leq\sqrt{d}\|v-v_{G(v)}\|_{\infty}\leq\sqrt{d}\tau_{1}\big{(}G(v)\big{)}\leq\sqrt{d}\tau_{1}(u)$. In the last inequality above, we used the definition of $u$ to deduce that $u\leq\tfrac{\mathcal{E}_{\infty}}{\sqrt{d}}.$
@@
(20) $\dfrac{\sqrt{d}e^{-\alpha\big{(}q-\tau_{3}(r)\big{)}}}{\rho_{T_{\alpha,\epsilon}}\big{(}B(v^{*},r)\big{)}}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big{(}\sqrt{2\mathcal{V}\big{(}\rho_{0}\big{)}}\Big{)}.$
@@
[…]
@@
(26) $\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\int_{\{G\geq u\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big{(}\rho_{0}\big{)}}\cdot\sqrt{\rho_{T_{\alpha,\epsilon}}\big{(}\{G\geq u\}\big{)}}\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big{(}\rho_{0}\big{)}}\cdot\dfrac{1}{\sqrt{u}}\cdot\sqrt{\int G\,d\rho_{T_{\alpha,\epsilon}}(v)}$
@@
where in the second inequality, we used the Cauchy inequality, in the third inequality, we used the Markov inequality and in the last inequality, we used the fact that $u\geq u(\delta).$ Thus by applying Lemma 4.10 with $B=C(0)$ and $\tilde{B}=\mathcal{V}(\rho_{0})$, when $\epsilon$ is small enough, the following holds:

### v8

@@ B. Assumptions on $G$:
Remark 3.
Assumption 2 (B1) is related to the convexity of function $G$ but is less stringent than the convexity condition. If it is not satisfied, similar to other gradient descent algorithms, there is a possibility for some particles to get trapped in the local minimizers $\hat{v}$ of $G$, i.e. $\nabla g(\hat{v})=0$. Nevertheless, provided the function values $\mathcal{E}(v)$ at those local minimizers of $G$ do not fall below $\mathcal{E}(v^{*})$, a condition attainable by adding a positive scalar multiple of $G$ to $\mathcal{E}$ without altering the solution $v^{*}$, it will not affect the convergence of the consensus point to the constrained minimizer $v^{*}$, as evidenced in the experiments detailed in Section 5.2.1, Figure 5. It is noteworthy that this slight adjustment on $\mathcal{E}(v)$ differs from the penalization method outlined in [8]. Here, there is no necessity for the penalty parameter to approach infinity, as the convergence is enforced through the dynamics rather than penalization. The introduction of a positive scalar multiple of $G$ to $\mathcal{E}$ is to avoid the extreme case. Thus a mild penalization would suffice.
Assumption 2 (B2) is primarily technical in nature. Assumption 2 (B3) guarantees that the gradient of $G$ around the constraint $\{G=0\}$ does not vanish too rapidly.
Although Assumption 2 (B) is formulated for an abstract function $G$, it is in fact satisfied by a rather large class of functions arising from geometric constraints. The next lemma shows that Assumption 2 (B) holds for the choice where $G$ is the squared distance to a compact convex constraint set $K$ of small intrinsic dimension. This covers in particular many constrained optimization problems in which the feasible set is contained in a lower–dimensional affine subspace, while projections onto $K$ remain computable, including common combinations of linear equality/inequality constraints with norm bounds.
A practical example is portfolio optimization in finance, where $v\in\mathbb{R}^{d}$ denotes portfolio weights (with $v_{j}$ the position in asset $j$), see [40]. Typical constraints include a budget constraint $\sum_{i=1}^{d}v_{i}+c=0$, market neutrality $\beta\cdot v=0$ (with $\beta$ the vector of asset betas), diversification $\|v\|_{2}^{2}\leq D$, and leverage $\|v\|_{1}\leq 2$. Intersecting such constraints yields a compact convex feasible set whose intrinsic dimension can be at most d2d-2. Common optimization objectives in this application include, for example, variance [35], $\beta$-Value-at-Risk [33, 36], and Conditional Value-at-Risk [45].
Lemma 4.2.
[Proof in Appendix B] Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set whose affine hull has dimension at most d2d-2. Define
$g(v):=\operatorname{dist}(v,K),\qquad G(v):=g(v)^{2},\qquad v\in\mathbb{R}^{d}.$
Then $G$ satisfies Assumption 2 (B).
@@
Remark 4.
The above Assumption 2 (C1) ensures the geometry of $\mathcal{E}$ on the set $\{G=u\}$ is similar among small enough $u$, i.e., on sets $\{G=u\}$, the constrained minimizers $v_{u}$ and constrained minimums $\mathcal{E}(v_{u})$ are close among small enough values for $u$. To illustrate, if this condition is not met, as depicted in Figure 3 (a), the desired constrained minimizer $v^{*}$ (depicted as a solid green pentagon) is considerably distant from the minimizer $v_{u}$ on a nearby level set $\{G=u\}$ (depicted as a solid orange circle) for all sufficiently small $u$. Consider an extreme case where we assume $\mathcal{E}(v_{u})$ is significantly smaller than $\mathcal{E}(v^{*})$ for all positive but sufficiently small $u$. Due to the nature of gradient descent on $G$ for each particle, which may not precisely enforce each particle to remain on the constraint $\{G=0\}$, these particles will tend to remain in a neighborhood of $\{G=0\}$. Consequently, numerous particles will cluster around $\{G=u\}$ for sufficiently small $u$, as illustrated in Figure 3 (a). Given that numerous particles are near $v_{u}$, where the function value is significantly small, the algorithm computes the consensus point around $v_{u}$ rather than $v^{*}$. Consequently, the consensus point will gradually lead particles to concentrate around $v_{u}$ rather than $v^{*}$, as illustrated in Figure 3 (b), which leads to a failure in this extreme scenario. To avoid the occurrence of such extreme cases, we proposed Assumption 2 (C1). In conjunction with other assumptions, the similarity of the local geometry of $\mathcal{E}$ on the set $\{G=u\}$ for sufficiently small values of $u$ is guaranteed as established in Lemma 4.4.
@@
Figure 3. The blue curves represent function values on the constraint set $\{G=0\}$ and the red curves on the level set $\{G=u\}$. Dashed lines represent corresponding constraint sets. The green pentagon denotes the constrained minimizer $v^{*}$, and the orange circle represents the minimizer $v_{u}$ on the nearby level set G=u{G=u}. Empty circles represent particles.
4.3. Sketch of the Proof
@@
We first plug $\tfrac{1}{2}\|v-v^{*}\|_{2}^{2}$ into Definition 3.2, which yields the following differential inequality that describes the dynamics of of the energy functional $\mathcal{V}\big(\rho_{t}\big)$.
@@
(Proof in Appendix E) Let $\mathcal{V}\big(\rho_{t}\big)$ be the energy functional defined in (9). Under Assumption 2,
(12) $\dfrac{d}{dt}\mathcal{V}\big(\rho_{t}\big)\leq$ $-(2\lambda-\sigma^{2})\mathcal{V}\big(\rho_{t}\big)+\sqrt{2}(\lambda+\sigma^{2})\sqrt{\mathcal{V}\big(\rho_{t}\big)}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$
$+\dfrac{\sigma^{2}}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}-\dfrac{1}{\epsilon}\int\Bigl\langle\nabla G,v-v^{*}\Bigr\rangle\,d\rho_{t}(v).\vskip-6.45831pt$
@@
to which Gronwall’s inequality can be applied, ensuring exponential decay. To this end, we first present the following auxiliary lemma, which draws out additional consequences from Assumption 2 (C).
@@
(Proof in Appendix C) There exist non-negative increasing function $\tau_{2}(x),\tau_{3}(x)$ and $\tau_{4}(x)$ mapping from $\mathbb{R}$ to $\mathbb{R}$ with $\lim_{x\rightarrow 0}\tau_{i}(x)=0$ for $i=2,3,4$ so that the following hold $\forall u,r\geq 0$ small enough:
@@
(Proof in Appendix G) Fix $r>0$ small enough and $u>0$ small enough. $q>0$ is a constant such that $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ is true $\forall\tilde{u}\in[0,u)$. Then
(14) $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}\leq$ $2\sqrt{d}\cdot\dfrac{\big(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\big)^{\mu}}{\eta}$
@@
It is observed that with an appropriate choice of $q,u,r$, provided $\rho_{t}(B(v^{*},r))$ is suitably bounded from below, as proven in Lemma 4.6 in Subsection 4.4, letting $\alpha$ be sufficiently large will make the first four terms above small enough. Concerning the last term of (12), it is related to $\int G\,d\rho_{t}(v)$, which can be controlled by Lemma 4.7 in Subsection 4.4. Consequently, we can control $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$ in such a way that (13) holds, thereby ensuring exponential decay.
Remark 5.
The framework of the proof strategy was introduced in [23]. However, in our case, the Laplace principle shown in [23] is not directly applicable. There are two challenges in our proof. First, a global valley-like structure around $v^{*}$ is required to use the Laplace principle. In the constrained case, such a structure does not hold. But one can notice that, on a nearby level set $\{G=u\}$, if a comparable constrained valley-like structure exists for the corresponding constrained minimizer $v_{u}$, a constrained Laplace principle on $\{G=u\}$ can be established for small $u$. Integrating over all $u$ yields reasonable estimates on $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|$. Secondly, there is always mass around the unconstrained global minimum. Consequently, increasing $\alpha$ would attract all particles towards the unconstrained global minimum rather than the constrained one. Therefore, one needs to find the balance between the gradient force and the tendency towards the consensus point.
@@
4.4. Proof of Theorem 4.1
In this subsection, we present the complete proof of Theorem 4.1. For simplicity, in the following, we assume $\tau_{1}(u)=\tau_{2}(u)=\tau_{3}(u)=\tau_{4}(u)=u$, where $\tau_{1},\ \tau_{2},\ \tau_{3},\ \tau_{4}$ are defined in Assumption 2 (C) and Lemma 4.4. We point out that the proof technique remains valid for any choice of $\tau_{i}$ that is an increasing function and converges to 0 as $u$ approaches 0.
@@
(Proof in Appendix H) Let $B=\sup_{t\in[0,T]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{\infty}$. Then for all $t\in[0,T]$,
@@
Dynamics of $\int G\,d\rho_{t}(v)$ In Lemma 4.3, we have gained control over $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$, yet the last term in the dynamics (12) remains to be studied, which we do now.
@@
(Proof in Appendix I) Assume $\sup_{t\in[0,T]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}<\infty$ and $\sup_{t\in[0,T]}\mathcal{V}\big(\rho_{t}\big)<\infty$. Then for $\epsilon>0$ small enough, $\int G\,d\rho_{t}(v)\leq\int G\,d\rho_{0}(v)$ $\forall t\in[0,T]$.
@@
Proof of Theorem 4.1.
First we use Lemma 4.3 to derive the dynamics of $\mathcal{V}\big(\rho_{t}\big)$:
$\dfrac{d}{dt}\mathcal{V}\big(\rho_{t}\big)\leq$ $-(2\lambda-\sigma^{2})\mathcal{V}\big(\rho_{t}\big)+\sqrt{2}(\lambda+\sigma^{2})\sqrt{\mathcal{V}\big(\rho_{t}\big)}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$
$+\dfrac{\sigma^{2}}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|^{2}_{2},$
where the last term on the right-hand-side of Lemma 4.3 is omitted because of its non-positivity due to Assumption 2 (B1).
Now we define $C(t)$ as $C(t)=\min\left\{\tfrac{\tau}{2}\tfrac{(2\lambda-\sigma^{2})}{\sqrt{2}(\lambda+\sigma^{2})},\sqrt{\tau\tfrac{(2\lambda-\sigma^{2})}{\sigma^{2}}}\right\}\sqrt{\mathcal{V}\big(\rho_{t}\big)}$ , and $T_{\alpha,\epsilon}$ as $T_{\alpha,\epsilon}=\sup\Big\{t\geq 0|\text{ }\mathcal{V}\big(\rho_{t^{\prime}}\big)>\delta,\big\|v_{\alpha}(\rho_{t^{\prime}}\big)-v^{*}\|_{2}\leq C(t^{\prime})\text{ for all $t^{\prime}\in[0,t]$}\Big\}.$ As long as $\|v_{\alpha}\big(\rho_{t^{\prime}}\big)-v^{*}\|_{2}\leq C(t^{\prime})$ is true, it is straightforward to verify that $\tfrac{d}{dt}\mathcal{V}\left(\rho_{t^{\prime}}\right)\leq-(1-\tau)(2\lambda-\sigma^{2})\mathcal{V}\left(\rho_{t^{\prime}}\right).$ Thus by Gronwall’s inequality, if $t\leq T_{\alpha,\epsilon}$, one has $\mathcal{V}\big(\rho_{t}\big)\leq\mathcal{V}(\rho_{0})\exp\left(-(1-\tau)(2\lambda-\sigma^{2})t\right).$
@@
One can verify that this choice of $q,r$ and $u$ will satisfy the assumptions of Lemma 4.5, i.e., $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ (for details, please see Appendix J).
@@
$\|v_{\alpha}\big(\rho_{T_{\alpha,\epsilon}}\big)-v^{*}\|_{2}\leq 2\sqrt{d}\cdot\dfrac{\big(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\big)^{\mu}}{\eta}$
@@
(17) $2\sqrt{d}\cdot\dfrac{\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$\leq 2\sqrt{d}\cdot\dfrac{\left(4\cdot\tfrac{1}{4}\cdot\left(\eta\tfrac{C\left(T_{\alpha,\epsilon}\right)}{8\sqrt{d}}\right)^{1/\mu}\right)^{\mu}}{\eta}=\dfrac{C\left(T_{\alpha,\epsilon}\right)}{4},$
@@
where $a(\delta)=2d\max\left\{\tfrac{\lambda\left(\sqrt{c}R_{0}+C(0)\right)\sqrt{c}}{(1-c)^{2}r(\delta)}+\tfrac{\sigma^{2}\left(cR_{0}^{2}+C(0)^{2}\right)(2c+1)}{(1-c)^{4}r(\delta)^{2}},\tfrac{2\lambda^{2}}{(2c-1)\sigma^{2}}\right\}$. In the first inequality above, we used (18), the fact that $T_{\alpha,\epsilon}<T^{*}$ and Lemma 4.6 with parameter $B=\sup_{t\in[0,T_{\alpha,\epsilon}]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{\infty}\leq\sup_{t\in[0,T_{\alpha,\epsilon}]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}\leq\sup_{t\in[0,T_{\alpha,\epsilon}]}C(t)\leq C(0)$. In the second inequality above, we used the Cauchy inequality. Also, Assumption 2 (C1) was used to deduce $\|v-v_{G(v)}\|_{2}\leq\sqrt{d}\|v-v_{G(v)}\|_{\infty}\leq\sqrt{d}\tau_{1}\big(G(v)\big)\leq\sqrt{d}\tau_{1}(u)$. In the last inequality above, we used the definition of $u$ to deduce that $u\leq\tfrac{\mathcal{E}_{\infty}}{\sqrt{d}}.$
@@
(20) $\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(r)\big)}}{\rho_{T_{\alpha,\epsilon}}\big(B(v^{*},r)\big)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\Big).$
@@
LHS of (23)
$\leq 2\cdot\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q(\delta)/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\mathcal{E}_{\infty}\Big)\leq\dfrac{1}{4}C_{\delta}\leq\dfrac{1}{4}C\big(T_{\alpha,\epsilon}\big),$
@@
(26) $\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\int_{\{G\geq u\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\sqrt{\rho_{T_{\alpha,\epsilon}}\big(\{G\geq u\}\big)}$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u}}\cdot\sqrt{\int G\,d\rho_{T_{\alpha,\epsilon}}(v)}$
@@
where in the second inequality, we used the Cauchy inequality, in the third inequality, we used the Markov inequality and in the last inequality, we used the fact that $u\geq u(\delta).$ Thus by applying Lemma 4.7 with $B=C(0)$ and $\tilde{B}=\mathcal{V}(\rho_{0})$, when $\epsilon$ is small enough (a quantification is given in Appendix I, Equation (73)), the following holds: